# Complexity of deep computations via topology of function spaces

### v2

@@ 4.1. NIP and Monte Carlo computability of deep computations
The raison d’être of the Baire class-1 functions is have a class of functions that contains the continuous functions and is closed under pointwise limits. By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$. This motivates the following definition:
@@
A function $f:X\rightarrow\mathbb{R}$ is universally measurable if and only if $f^{-1}(U)$ is $\mu$-measurable for every Radon probability measure $\mu$ on $X$ and every open set $U\subseteq\mathbb{R}$.
@@
In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the set $\pi_{i}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open). Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
@@
    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has an accumulation point in $\mathscr{M}^{0}(X,\mu)$.
@@
    Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”. The formal statement of this, which is Luzin’s theorem, is that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ is a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable ) equipped with a Borel algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|_{F}$ is continuous and $\mu(E\backslash F)<\varepsilon$.
@@
Condition (2) of Theorem  shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$. Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure. Therefore, in the context of Monte Carlo measurability, we focus on Radon probability measures rather than general Radon measures.
@@ 4.2. Talagrand stability and Monte Carlo computability of deep computations
There is another notion closely related to NIP, introduced by Talagrand in [TAL84] while studying Pettis integration. Suppose that $X$ is a compact Hausdorff space and $A\subseteq\mathbb{R}^{X}$. Let $\mu$ be a Radon probability measure on $X$. Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$. we write:
@@
We say that $A$ is Talagrand $\mu$-stable if and only if for every $\mu$-measurable set $E\subseteq X$ of positive measure and for every $a<b$ there is $k\geq 1$ such that
@@
where $\mu^{*}$ denotes the outer measure (we need to work with outer measure since the sets $D_{k}(A,E,a,b)$ need not be $\mu$-measurable). The inequality certainly holds when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
The main result of this section is that accumulation points of standard computations (i.e., deep computations) are Monte Carlo computable; this is Theorem 4.14 below. We now prove that deep computations that are Talagrand $\mu$-stable are $\mu$-measurable. But first, let us state the following useful characterization of measurable functions (compare with Fact 1.2):
@@
(In particular, if $X$ is Hausdorff, $\mu$ is a Radon measure on $X$ and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold). Then, $f:X\rightarrow\mathbb{R}$ is measurable if and only if for every $K\in\mathcal{K}$ with $0<\mu(K)<\infty$ and $a<b$, either $\mu^{*}(P)<\mu(K)$ or $\mu^{*}(Q)<\mu(K)$ where $P=\{x\in K:f(x)\leq a\}$ and $Q=\{x\in K:f(x)\geq b\}$.
The following technical lemma will be instrumental to prove of Proposition 4.13, which, in turn, will yield the main result of the subsection, Theorem 4.14.
@@
We first observe that universal Talagrand stability corresponds to a complexity class smaller than or equal to the NIP:
@@
On the other hand, by fixing a particular well-known measure, the Lebesgue measure, we see that the other case is also consistent:
@@
Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions. It is possible that the continuity assumption, i.e., the Extendibility Axiom, settles the problem. This is a problem that remains open for future work.

### v3

@@ 4.1. NIP and Monte Carlo computability of deep computations
The raison d’être of the Baire class-1 functions is to have with a class of functions that are obtained as equential limit points of continuous functions. By Fact 1.2, for perfectly normal $X$, a function $f$ is in $B_{1}(X,Y)$ if and only if $f^{-1}[U]$ is an $F_{\sigma}$ subset of $X$ for every open $U\subseteq Y$. Thus, for such $X$, functions in $B_{1}(X,Y)$ are not too far from being continuous. In this section we will study a more general class of functions, namely, the class of universally measurable functions, which we define next.
@@
If $X$ is a compact (Hausdorff) space, then every Radon measure $\mu$ on $X$ is finite. Then, the measure given by $\nu(A):=\mu(A)/\mu(X)$ is a probability measure on $X$ with the same null sets as $\mu$. Hence, Radon measures on compact spaces are equivalent to (Radon) probability measures. We summarize this fact in the next remark:
@@
If $X$ is compact, then a function $f:X\rightarrow\mathbb{R}$ is universally measurable if and only if $f^{-1}(U)$ is $\mu$-measurable for every Radon probability measure $\mu$ on $X$ and every open set $U\subseteq\mathbb{R}$.
@@
In the context of deep computations, we are interested in transition maps of a state space $L\subseteq\mathbb{R}^{\mathcal{P}}$ into itself. There are two natural $\sigma$-algebras one can consider in the product space $\mathbb{R}^{\mathcal{P}}$: the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by open sets in $\mathbb{R}^{\mathcal{P}}$, and the cylinder $\sigma$-algebra (i.e., the $\sigma$-algebra generated by the sub-basic open sets in $\mathbb{R}^{\mathcal{P}}$, that is, the sets $\pi_{P}^{-1}(U)$ with $U\subseteq\mathbb{R}$ open and $P\in\mathcal{P}$). Note that when $\mathcal{P}$ is countable, both $\sigma$-algebras coincide, but in general the cylinder $\sigma$-algebra is strictly smaller. We will use the cylinder $\sigma$-algebra to define universally measurable maps $f:\mathbb{R}^{\mathcal{P}}\rightarrow\mathbb{R}^{\mathcal{P}}$. The reason for this choice is the following characterization:
@@
    For every Radon measure $\mu$ on $X$, AA is relatively countably compact in $\mathscr{M}^{0}(X,\mu)$, i.e., every countable subset of $A$ has a limit point in $\mathscr{M}^{0}(X,\mu)$.
@@
    Littlewoood’s second principle states that every Lebesgue measurable function is “nearly continuous”. The formal statement of this, which is Luzin’s theorem, is that if $(X,\Sigma,\mu)$ a Radon measure space and $Y$ is a second-countable topological space (e.g., $Y=\mathbb{R}^{\mathcal{P}}$ with $\mathcal{P}$ countable) equipped with the Borel $\sigma$-algebra, then any given $f:X\to Y$ is measurable if and only if for every $E\in\Sigma$ and every $\varepsilon>0$ there exists a closed $F\subseteq E$ such that the restriction $f|_{F}$ is continuous and $\mu(E\backslash F)<\varepsilon$.
@@
Condition (2) of Theorem 4.5 shows that to study measure-theoretic versions of NIP, we need only consider compact subsets of $X$. Now, every Radon measure on a compact space is finite; hence, by proper normalization, it can be treated as a probability measure. Therefore, in the context of Monte Carlo measurability, we focus on Radon probability measures rather than general Radon measures.
@@ 4.2. Talagrand stability and Monte Carlo computability of deep computations
There is another notion closely related to NIP, introduced by Talagrand in [TAL84] while studying Pettis integration. Suppose that $X$ is a compact Hausdorff space and $A\subseteq\mathbb{R}^{X}$. Let $\mu$ be a Radon probability measure on $X$. Given a $\mu$-measurable set $E\subseteq X$, a positive integer $k$ and real numbers $a<b$, we write:
@@
We say that $A$ is Talagrand $\mu$-stable if and only if for every $\mu$-measurable set $E\subseteq X$ of positive measure and for every $a<b$ there is a $k\geq 1$ such that
@@
where $\mu^{*}$ denotes the outer measure (we need to work with outer measure since the sets $D_{k}(A,E,a,b)$ need not be $\mu^{2k}$-measurable). The inequality certainly holds when $A$ is a countable set of continuous (or $\mu$-measurable) functions.
The main result of this section is that deep computations, i.e., limit points of a Talagrand stable set of computations are Monte Carlo computable; this is Theorem 4.14 below. We now prove that limit points of a Talagrand $\mu$-stable set are $\mu$-measurable. But first, let us state the following useful characterization of measurable functions (compare with Fact 1.2):
@@
(In particular, if $X$ is compact Hausdorff, $\mu$ is a Radon probability measure on $X$, $\Sigma$ is the completion of the Borel $\sigma$-algebra by $\mu$, and $\mathcal{K}$ is the collection of compact subsets of $X$, all these conditions hold). Then, $f:X\rightarrow\mathbb{R}$ is measurable if and only if for every $K\in\mathcal{K}$ with $0<\mu(K)<\infty$ and $a<b$, either $\mu^{*}(P)<\mu(K)$ or $\mu^{*}(Q)<\mu(K)$ where $P=\{x\in K:f(x)\leq a\}$ and $Q=\{x\in K:f(x)\geq b\}$.
The following technical lemma will be instrumental for proving Proposition 4.13, which, in turn, will yield the main result of the subsection, namely Theorem 4.14.
@@
We first observe that universal Talagrand stability corresponds to a complexity class smaller than or equal to the NIP class:
@@
On the other hand, by fixing a particular well-known measure, namely the Lebesgue measure, we see that the other case is also consistent:
@@
Notice that the preceding two results apply to sets of measurable functions, a class of functions larger than the class of continuous functions. However, by the Extendibility Axiom, finitary computations are continuous, i.e., if $A$ is a set of computations, then $A\subseteq C_{p}(X)$. The question of whether we can remove the set-theoretic assumption in Theorem 4.15 when $A\subseteq C_{p}(X)$ (instead of $A\subseteq M_{r}(X)$) remains open.